OpenStudy (anonymous):

3x+6y=90 what is y intercept and x intercept? how are you getting the answer? I am having trouble with this...

OpenStudy (anonymous):

first put x=0 and u get 3(0) + 6y = 90 or 6y = 90 or y = 90/6 = 15 this is the y intercept similarly put y=0 and calculate to get the x-intercept...

OpenStudy (anonymous):

for y intercept put x=0 and for x intercept put y=0 and ur answers are 15 and 30 respectively

OpenStudy (anonymous):

for y=0 3x + 6(0) = 90 3x = 90 x= 90/3 x= 30 this is your x intercept u can use this for all other questions for y intercept put x=0 and calculate for x intercept put y=0 and calculate

OpenStudy (anonymous):

I think I understand it a little better, now thank you

OpenStudy (anonymous):

u r welcome :D